OpenStudy (mhchen):

Brief explanation of how intervals work, and what they measure? Opportunity for a shinyyy medal as well :)

OpenStudy (mathmale):

Please ask this question in context. What are you using intervals for? What does "work" mean in terms of "how intervals work?"

OpenStudy (studygurl14):

What do you mean by "intervals"?

OpenStudy (mathmale):

@mhchen?

OpenStudy (mhchen):

You know, the giant S thing

OpenStudy (studygurl14):

Integral?

OpenStudy (mhchen):

Woops typo

OpenStudy (studygurl14):

Do you have a specific integral question?

OpenStudy (mhchen):

Nope, I just want to know how it's used.

OpenStudy (mathmale):

No, I don't know "the giant S thing." It is essential that you learn the correct vocabulary, or you'll lose a lot of time, and so will others who are trying to help you. The topic of "integrals" is far too broad to give you a simple, direct answer. Please be specific about what you need to learn.

OpenStudy (mhchen):

Geez mathmale...

OpenStudy (studygurl14):

An integral, very simply put, is the area bounded by a graph. But beyond that, I really need a specific question to explain further.

OpenStudy (mhchen):

Alrighty then.

OpenStudy (mathmale):

Sorry my insistence on correctness bothers you. An integral is the opposite of a derivative. If you start with function f(x), take the derivative f '(x), and then integrate your result, you'll arrive back at your original function, "plus C." C is the constant of integration.

OpenStudy (studygurl14):

That too^ See @mhchen there are many applications of integrals.

OpenStudy (mhchen):

Aaaah, so it's the opposite of derivatives, that makes more sense.

OpenStudy (mathmale):

Example: the derivative of f(x)=x^3 is f '(x)=3x^2. Now, if you were to integrate:\[\int\limits_{}^{} 3x^2 dx\] the correct integrand would be x^3+C.